How would I set this problem up? A Cupronickel alloy that is 60% copper and 40% nickel is needed to manufacturea marine propeller. Only two cupronickel alloys are available, one is 75% copper and 25% nickel and the other is 55% copper and 45% nickel. How many pounds of each of these alloys need to be melted and combined to obtain 300 pounds of the 60% copper/ 40% nickel alloy? I am havin trouble because there are two metals that need to be melted together and I'm not sure of what I am supposed to do.
Answer by mananth(16946)

The problem is set to just see your analatical ability. ( OR confuse you)
If you have 60% copper then naturally you will have 40% Nickel in the final alloy.
So lets say we need 60% copper in the alloy.
--------percent ----------------quantity
Alloy 1 75 ----------------x pounds
Alloy 2 55 ----------------300-x pounds
Total 60.00% ----------------300
75x+55(300-x)= 60*300
75x+16500-55x=18000
75x-55x =18000 -16500
20x=1500
/20
x=75 pounds 75.00% Alloy
225 pounds of 55.00% Alloy
---------------------
Work the other way using Nickel as the required %. you get the same quantities.
---------percent----------------quantity
Alloy 1 25 ----------------x pounds
Alloy 2 45 ----------------300-x pounds
Total 5.00% ----------------300
25x+45(300-x)=40*300
25x+13500-45x=12000
25x-45x =12000-13500
-20x= -1500
/-20
x=75 pounds 25.00%


225 pounds 45.00%